can you please solve this below problem?
Mrs. Rifkin had 84 oranges. 24 of her oranges were rotten, and the rest were either ripe or unripe. There were 18 more ripe oranges than unripe oranges. Find the ratio of the number of ripe oranges to the number of rotten oranges to the number of unripe oranges?

Answer by stanbon(75887)

Mrs. Rifkin had 84 oranges. 24 of her oranges were rotten, and the rest were either ripe or unripe. There were 18 more ripe oranges than unripe oranges. Find the ratio of the number of ripe oranges to the number of rotten oranges to the number of unripe oranges?
-----------------------
rotten = 24
----------------------------------
ripe + unripe = 84-24 = 60
ripe - unripe = 18
-----------------------
2*ripe = 78
ripe = 39
----
Ans:: ripe/rotten = 39/24 = 13/8
------------------
Cheers,
Stan H.
------------